# Improving Time Series Forecasting via Instance-aware Post-hoc Revision

### v1

@@ Abstract
Time series forecasting plays a vital role in various real-world applications and has attracted significant attention in recent decades. While recent methods have achieved remarkable accuracy by incorporating advanced inductive biases and training strategies, we observe that instance-level variations remain a significant challenge. These variations—stemming from distribution shifts, missing data, and long-tail patterns—often lead to suboptimal forecasts for specific instances, even when overall performance appears strong. To address this issue, we propose a model-agnostic framework, PIR, designed to enhance forecasting performance through Post-forecasting Identification and Revision. Specifically, PIR first identifies biased forecasting instances by estimating their accuracy. Based on this, the framework revises the forecasts using contextual information, including covariates and historical time series, from both local and global perspectives in a post-processing fashion. Extensive experiments on real-world datasets with mainstream forecasting models demonstrate that PIR effectively mitigates instance-level errors and significantly improves forecasting reliability.
@@ 1 Introduction
Time series forecasting is a fundamental task in time series analysis, attracting considerable attention in recent years [39, 47]. Various applications have been facilitated by the advancement of forecasting, including traffic planning [17], stock market prediction [23], healthcare analytics [19], and weather forecasting [4, 57]. Recent years have witnessed significant efforts dedicated to this area, with deep learning-based approaches achieving remarkable success due to their powerful ability to capture both temporal [60, 61] and cross-channel dependencies [58, 30]. Furthermore, advanced inductive biases and training strategies have been introduced to address the non-stationary nature of time series data [20, 33] and to construct foundation models for time series forecasting [10, 32, 50].
Despite their satisfactory performance in overall evaluations, we emphasize that existing forecasting approaches often overlook inherent instance-level variations, which can arise from the long-tail distribution of numerical patterns in time series data and ultimately lead to forecasting failures in specific cases. Specifically, time series typically represent the numerical reflections of complex and dynamic real-world systems and are therefore prone to noise, sensor failures, and other anomalies during data collection. These issues can result in potential distribution shifts [26], missing values [41], or unforeseen anomalies [5]. Therefore, while most instances exhibit similar numerical patterns, a subset may present rare behaviors, and mainstream forecasting methods often struggle to effectively model these exceptional instances, leading to inaccurate or even unreliable forecasting outcomes.
@@
To better illustrate this phenomenon, we present the per-instance Mean Squared Error (MSE) evaluation curve of PatchTST [36] on the ETTh1 dataset [60], along with the corresponding error distribution visualized through both a histogram and a kernel density estimate in Figure 1. As shown in the figure, while the MSE remains consistently low for the majority of instances, there exist specific cases where PatchTST yields unsatisfactory forecasting results, as indicated by multiple spikes in the error curve. Moreover, the error distribution clearly exhibits a long-tail pattern, reinforcing our motivation and underscoring the challenges posed by the instance-level variations.
To this end, we propose the PIR framework, designed to enhance forecasting results via Post-forecasting Identification and Revision, addressing the challenge from a novel post-processing perspective. The framework comprises two key components. The first is the failure identification mechanism, which identifies the potential biased forecasting instances where the model’s predictions are less reliable, through estimating the forecasting performance on a per-instance level. The second is a post-revision module, which refines the forecasts by leveraging contextual information from both local and global perspectives. For the local revision, inspired by exogenous variable modeling approaches [37, 48], we utilize immediate forecasts of covariates along with available exogenous information such as timestamps [45] and textual descriptions [27] as side information to implicitly mitigate the impact of instance-level variations within a local window. This approach is grounded in the assumption that lead-lag effects between covariates can provide valuable insights into future trends [59], and exogenous information can serve as additional prior conditions guiding the revision. For the global revision, the framework addresses rare or atypical numerical patterns by explicitly retrieving similar instances from the global historical data [28]. This retrieval-based strategy enables the model to better capture long-tail patterns that may be overlooked by conventional forecasting models. Finally, PIR integrates the original forecasts with its revision outputs via a weighted sum, making it model-agnostic and broadly compatible with existing forecasting architectures. The primary contributions of our work are summarized as follows:
@@ 2.1 Time Series Forecasting
Time series forecasting has been a central area of research for several decades. One of the earliest landmark contributions was the development of statistical methods, which are celebrated for their solid theoretical foundation and systematic approach to model design. Representative works in this domain include ARIMA and Holt-Winters [6, 18], which have laid the groundwork for more advanced methodologies. More recently, with the advancement of deep learning, numerous neural forecasting models have been proposed, achieving superior performance. Recurrent Neural Networks (RNNs) are among the first deep learning architectures applied to forecasting [49, 40], followed by a variety of other structures, such as convolution-based models [21, 35, 8], attention-based networks [22, 24, 61], and MLP-based architectures [38, 55, 25]. These models have demonstrated remarkable capabilities in capturing both temporal and cross-channel dependencies within time series data, leading to substantial improvements in forecasting accuracy.
Beyond advancements in model architecture, a range of specialized techniques rooted in time series analysis has also emerged. These include methods for time series decomposition [52, 54], frequency modeling [53], and approaches for non-stationary forecasting [20, 33], which address the unique challenges presented by time series data. Besides, self-supervised pretraining has gained significant attention as a powerful approach, with various training strategies being explored [51, 7, 56]. Moreover, recent efforts aim to construct foundational time series models capable of learning universal representations and being applied to diverse downstream tasks [34, 10, 32, 50].
@@ 2.2 Context Modeling in Time Series Forecasting
In addition to forecasting based solely on multivariate time series data, several pioneering studies have explored the incorporation of contextual information to enhance forecasting performance [2]. On one hand, some advancements focus on integrating exogenous information within a local window during the forecasting process. For instance, TFT [24] and TiDE [9] leverage dense encoders to process timestamp information, which is subsequently used to condition future forecastings. Similarly, NBEATSx [37] and TimeXer [48] improves the forecasting accuracy of target variables by explicitly modeling the influence of exogenous variables. On the other hand, some studies investigate the potential of retrieving relevant time series from the global historical context. RATD [28] utilizes the retrieved series as references to guide the denoising process of diffusion-based forecasters, and RAF [43] introduces a principled retrieval-augmented generation framework for time-series forecasting. These approaches both demonstrate promising results.
@@ 3.2 Failure Identification
The goal of the PIR framework is to post-process forecasting results by identifying and revising potential forecasting failures caused by instance-level variations. Thus, the identification process becomes a primary target and challenge. Conceptually, it forms an Uncertainty Estimation task based on the immediate forecasting results. Though various efforts have been devoted in this area, it has not been deeply explored for properly estimating the uncertainty of point-wise forecasting results mainly due to two reasons. Firstly, given the regression nature of the forecasting task, most existing approaches generate predictions directly from the hidden states. As a result, token-level distributions before generation are not available, and common uncertainty evaluation methods based on probabilities are not directly applicable [3, 12]. Secondly, the forecasting failures arise from both data uncertainty (e.g., missing values) and model uncertainty (e.g., underfitting on long-tail patterns) [13] while most existing uncertainty quantification methods primarily focus on the latter one.
To this end, we innovatively quantify uncertainty using forecasting error and adopt a data-driven approach as a practical solution for estimating the uncertainty of a given input-forecast pair $(x,\bar{y})$. Specifically, following common practice in modeling complex dynamics [31, 11], we employ a two-layer fully connected neural network with non-linear activation functions $f_{ue}(\cdot)$ to estimate the forecasting uncertainty. Additionally, we introduce an auxiliary constraint to support the abovementioned procedure. Since explicit uncertainty is not directly available for the forecasting results, we treat the forecasting error as a feasible guiding signal. The intuition behind this is that higher uncertainty will likely result in greater forecasting error. Therefore, the estimated uncertainty is further constrained to predict the MSE of the forecasting results:
@@ 3.3 Local Revising
After identifying the potential forecasting failures with high uncertainty, the PIR framework revises these results from both local and global perspectives. For the local revision component, the core idea is to leverage the contextual information within a local horizon window, which includes the intermediate forecastings of covariates and available exogenous variables, to enhance forecasting accuracy. The intuition behind this approach is straightforward. Firstly, the lead-lag effects are widely present in the time series data, where the forecasting results of covariates can provide valuable insights into future trends [59]. Besides, the exogenous variables, such as the time-related information and environmental factors known in advance can serve as a prior condition [24, 45], mitigating the impact of sudden distribution shifts caused by natural rhythms like holiday effects [15, 42]. Conceptually, this approach is particularly beneficial for models that prioritize robustness over capacity by adopting a channel-independent strategy [14].
In practice, we project the intermediate forecasting results into hidden states on a per-variate basis [30], along with the available exogenous variable information, which are concatenated for further correlation extraction. Let $h_{co},h_{exo}$ be the representation of covariates and exogenous variables respectively, and $c$ refer to the available exogenous variable information corresponding to the instance, the projection process is formulated as:
@@
Here $\text{CoVariateEmb}(\cdot)$ is a trainable linear projector, and $\text{ExoVariateEmb}(\cdot)$ is implemented flexibly based on the characteristic of $c$, which can be a linear projector for numerical features or language models for textual descriptions [27]. Through this approach, local context from multiple sources and domains can be seamlessly embedded in the framework to guide the revising process.
The hidden states are then passed through a traditional Transformer [44] with a linear prediction head to generate the revised results $y_{local}$. By leveraging the attention mechanism, the local revision component explicitly captures the correlations between covariates and exogenous variables, thereby ensuring that the contextual information within the local window is fully utilized to correct forecasting failures and enhance prediction accuracy.
@@ 3.4 Global Revising
$\text{Index},w=$ TopK $\text{TopK }\operatorname{Sim}(Enc(x),Enc(X)),$ (3)
$Y_{re}$ =$=\{Y_{train,i}|\forall i\in\text{Index}\}.$
Here, $Enc(\cdot)$ is an encoding function that processes the time series, which is instantiated as an instance normalization operation [20] in practice for its simplicity and effectiveness of mitigating the impact of non-stationarity that could lead to unstable similarity estimation. Moreover, powerful pre-trained forecasting models [10, 32, 50] can be optionally utilized for better representation projection. Additionally, $w$ represents the top-$K$ similarity scores estimated by the similarity operator $\operatorname{Sim}(\cdot,\cdot)$, which is instantiated using cosine similarity due to its computational efficiency.
@@ 3.5 Optimization Target
By combining the components described above, the PIR framework produces the final forecasting results through post-forecasting identification and revision using a residual approach [16]. This process is expressed as:
@@
Here, $\alpha$ and $\beta$ are learned weights corresponding to the local and global revision components, respectively, and $\sigma(\cdot)$ denotes the Sigmoid activation function. A linear transformation is employed to estimate $\alpha$, with its weight and bias initialized to vectors of ones and zeros. This design ensures that higher uncertainty estimates lead to larger values of $\alpha$, thus placing greater emphasis on the local revision. In contrast, $\beta$ is generated through a multi-layer perceptron (MLP), which takes both the estimated uncertainty and the retrieval similarities as input. This allows the model to dynamically adjust the influence of the global revision based on the confidence of the prediction and the relevance of the retrieved historical series. The overall optimization objective is to minimize the mean squared error (MSE) between the revised forecasts and the ground truth, formulated as:
@@ 4.1 Experimental Setup
Datasets. For the long-term forecasting task, we conduct experiments on a widely recognized benchmark dataset that includes eight real-world datasets spanning diverse domains [52, 21]. Additionally, we incorporate the PEMS dataset, which contains four subsets, for the short-term forecasting task [29]. The exogenous information used in these datasets are the available timestamps. We also conduct experiments on datasets with additional textual descriptions in the Appendix. Following standard experimental protocols, we split each dataset into training, validation, and testing sets in chronological order. The split ratios are set to 6:2:2 for the ETT dataset and 7:1:2 for the other datasets. Detailed information about the datasets is available in the Appendix.
Backbone Models. PIR is a model-agnostic plugin that can be seamlessly integrated with any time series forecasting model. To demonstrate the effectiveness of the framework, we select four mainstream forecasting models based on diverse architectures, encompassing both channel-dependent and channel-independent assumptions, as backbones. These models are evaluated under both long-term and short-term forecasting settings: PatchTST [36], SparseTSF [25], iTransformer [30] and TimeMixer [46]. We implement these models following their official code.

### v2

@@ Abstract
Time series forecasting plays a vital role in various real-world applications and has attracted significant attention in recent decades. While recent methods have achieved remarkable accuracy by incorporating advanced inductive biases and training strategies, we observe that instance-level variations remain a significant challenge. These variations—stemming from distribution shifts, missing data, and long-tail patterns—often lead to suboptimal forecasts for specific instances, even when overall performance appears strong. To address this issue, we propose a model-agnostic framework, PIR, designed to enhance forecasting performance through Post-forecasting Identification and Revision. Specifically, PIR first identifies biased forecasting instances by estimating their accuracy. Based on this, the framework revises the forecasts using contextual information, including covariates and historical time series, from both local and global perspectives in a post-processing fashion. Extensive experiments on real-world datasets with mainstream forecasting models demonstrate that PIR effectively mitigates instance-level errors and significantly improves forecasting reliability. Our code is available [^1] https://github.com/icantnamemyself/PIR
@@ 1 Introduction
Time series forecasting is a fundamental task in time series analysis, attracting considerable attention in recent years [qiu2024tfb, wang2024deep]. Various applications have been facilitated by the advancement of forecasting, including traffic planning [jin2023spatio], stock market prediction [li2024master], healthcare analytics [kaushik2020ai], and weather forecasting [nature, nature2]. Recent years have witnessed significant efforts dedicated to this area, with deep learning-based approaches achieving remarkable success due to their powerful ability to capture both temporal [Informer, fedformer] and cross-channel dependencies [crossformer, itransformer]. Furthermore, advanced inductive biases and training strategies have been introduced to address the non-stationary nature of time series data [revin, san] and to construct foundation models for time series forecasting [timesfm, timer, moirai].
Despite their satisfactory performance in overall evaluations, we emphasize that existing forecasting approaches often overlook inherent instance-level variations, which can arise from the long-tail distribution of numerical patterns in time series data and ultimately lead to forecasting failures in specific cases. Specifically, time series typically represent the numerical reflections of complex and dynamic real-world systems and are therefore prone to noise, sensor failures, and other anomalies during data collection. These issues can result in potential distribution shifts [liu2024timeseries], missing values [tashiro2021csdi], or unforeseen anomalies [blazquez2021review]. Therefore, while most instances exhibit similar numerical patterns, a subset may present rare behaviors, and mainstream forecasting methods often struggle to effectively model these exceptional instances, leading to inaccurate or even unreliable forecasting outcomes.
@@
To better illustrate this phenomenon, we present the per-instance Mean Squared Error (MSE) evaluation curve of PatchTST [patchtst] on the ETTh1 dataset [Informer], along with the corresponding error distribution visualized through both a histogram and a kernel density estimate in Figure 1. As shown in the figure, while the MSE remains consistently low for the majority of instances, there exist specific cases where PatchTST yields unsatisfactory forecasting results, as indicated by multiple spikes in the error curve. Moreover, the error distribution clearly exhibits a long-tail pattern, reinforcing our motivation and underscoring the challenges posed by the instance-level variations.
To this end, we propose the PIR framework, designed to enhance forecasting results via Post-forecasting Identification and Revision, addressing the challenge from a novel post-processing perspective. The framework comprises two key components. The first is the failure identification mechanism, which identifies the potential biased forecasting instances where the model’s predictions are less reliable, through estimating the forecasting performance on a per-instance level. The second is a post-revision module, which refines the forecasts by leveraging contextual information from both local and global perspectives. For the local revision, inspired by exogenous variable modeling approaches [nbeatsx, timexer], we utilize immediate forecasts of covariates along with available exogenous information such as timestamps [wang2024rethinking] and textual descriptions [timemmd] as side information to implicitly mitigate the impact of instance-level variations within a local window. This approach is grounded in the assumption that dependencies between covariates like lead-lag effects can provide valuable insights into future trends [zhao2024rethinking], and exogenous information can serve as additional prior conditions guiding the revision. For the global revision, the framework addresses rare or atypical numerical patterns by explicitly retrieving similar instances from the global historical data [liu2024retrievalaugmented]. This retrieval-based strategy enables the model to better capture long-tail patterns that may be overlooked by conventional forecasting models. Finally, PIR integrates the original forecasts with its revision outputs via a weighted sum, making it model-agnostic and broadly compatible with existing forecasting architectures. The primary contributions of our work are summarized as follows:
@@ 2.1 Time Series Forecasting
Time series forecasting has been a central area of research for several decades. One of the earliest landmark contributions was the development of statistical methods, which are celebrated for their solid theoretical foundation and systematic approach to model design. Representative works in this domain include ARIMA and Holt-Winters [arima, holt-winters], which have laid the groundwork for more advanced methodologies. More recently, with the advancement of deep learning, numerous neural forecasting models have been proposed, achieving superior performance. Recurrent Neural Networks (RNNs) are among the first deep learning architectures applied to forecasting [wen2017multi, deepar], followed by a variety of other structures, such as convolution-based models [lstnet, moderntcn, convtimenet], attention-based networks [logtrans, tft, fedformer], and MLP-based architectures [nbeats, DLinear, sparsetsf]. These models have demonstrated remarkable capabilities in capturing both temporal and cross-channel dependencies within time series data, leading to substantial improvements in forecasting accuracy.
Beyond advancements in model architecture, a range of specialized techniques rooted in time series analysis has also emerged. These include methods for time series decomposition [wu2021autoformer, yu2024revitalizing], frequency modeling [yi2024frequency], and approaches for non-stationary forecasting [revin, san], which address the unique challenges presented by time series data. Besides, self-supervised pretraining has gained significant attention as a powerful approach, with various training strategies being explored [woo2022cost, timemae, tpami_ssl]. In addition, recent efforts aim to construct foundational time series models capable of learning universal representations and being applied to diverse downstream tasks [gpht, timesfm, timer, moirai]. Moreover, some pioneer works explore forecasting enhanced with the reasoning ability of LLMs [wang2025can, luo2025time].
@@ 2.2 Context Modeling in Time Series Forecasting
In addition to forecasting based solely on multivariate time series data, several pioneering studies have explored the incorporation of contextual information to enhance forecasting performance [ashokcontext]. On one hand, some advancements focus on integrating exogenous information within a local window during the forecasting process. For instance, TFT [tft] and TiDE [tide] leverage dense encoders to process timestamp information, which is subsequently used to condition future forecastings. Similarly, NBEATSx [nbeatsx] and TimeXer [timexer] improves the forecasting accuracy of target variables by explicitly modeling the influence of exogenous variables. On the other hand, some studies investigate the potential of retrieving relevant time series from the global historical context. RATD [liu2024retrievalaugmented] utilizes the retrieved series as references to guide the denoising process of diffusion-based forecasters, and RATSF [wang2024ratsf] introduces the retrieval augmented cross-attention architecture for explicitly modeling similar historical data. More recently, RAFT [raft] achieves satisfactory performance through forecasting enhanced with multi-period retrieval.
@@ 3.2 Failure Identification
The goal of the PIR framework is to post-process forecasting results by identifying and revising potential forecasting failures caused by instance-level variations. Thus, the identification process becomes a primary target and challenge. Conceptually, it forms an Uncertainty Estimation task based on the immediate forecasting results. Though various efforts have been devoted in this area, it has not been deeply explored for properly estimating the uncertainty of point-wise forecasting results mainly due to two reasons. Firstly, given the regression nature of the forecasting task, most existing approaches generate predictions directly from the hidden states. As a result, token-level distributions before generation are not available, and common uncertainty evaluation methods based on probabilities are not directly applicable [perplexity, gal2016uncertainty]. Secondly, the forecasting failures arise from both data uncertainty (e.g., missing values) and model uncertainty (e.g., underfitting on long-tail patterns) [gawlikowski2023survey] while most existing uncertainty quantification methods primarily focus on the latter one.
Although we have identified several potential reasons that may lead to prediction failure, there are many other potential and interrelated factors, and there is a lack of ground truth and metrics to identify and disentangle the specific reasons for each forecasting failure instance. Therefore, we do not explicitly locate these failure patterns. Instead, we innovatively quantify uncertainty using forecasting error and adopt a data-driven approach as a practical solution for estimating the uncertainty of a given input-forecast pair $(x,\bar{y})$, which is more flexible and can potentially accommodate a broader range of failure modes. Specifically, following common practice in modeling complex dynamics [nst, fan2023dish], we employ a two-layer fully connected neural network with non-linear activation functions $f_{ue}(\cdot)$ to estimate the forecasting uncertainty. Additionally, we introduce an auxiliary constraint to support the abovementioned procedure. Since explicit uncertainty is not directly available for the forecasting results, we treat the forecasting error as a feasible guiding signal. The intuition behind this is that higher uncertainty will likely result in greater forecasting error. Therefore, the estimated uncertainty is further constrained to predict the MSE of the forecasting results:
@@ 3.3 Local Revising
After identifying the potential forecasting failures with high uncertainty, the PIR framework revises these results from both local and global perspectives. For the local revision component, the core idea is to leverage the contextual information within a local horizon window, which includes the intermediate forecastings of covariates and available exogenous variables, to enhance forecasting accuracy. The intuition behind this approach is straightforward. Firstly, the dependencies including lead-lag effects are widely present in the time series data, where the forecasting results of covariates can provide valuable insights into future trends [zhao2024rethinking]. Besides, the exogenous variables, such as the time-related information and environmental factors known in advance can serve as a prior condition [tft, wang2024rethinking], mitigating the impact of sudden distribution shifts caused by natural rhythms like holiday effects [harvey1990estimation, taylor2018forecasting]. Conceptually, this approach is particularly beneficial for models that prioritize robustness over capacity by adopting a channel-independent strategy [han2024capacity].
In practice, we project the intermediate forecasting results into hidden states on a per-variate basis [itransformer], along with the available exogenous variable information, which are concatenated for further correlation extraction. Let $h_{co},h_{exo}$ be the representation of covariates and exogenous variables respectively, and $c$ refer to the available exogenous variable information corresponding to the instance, the projection process is formulated as:
@@
Here $\text{CoVariateEmb}(\cdot)$ is a trainable linear projector, and $\text{ExoVariateEmb}(\cdot)$ is implemented flexibly based on the characteristic of $c$, which can be a linear projector for numerical features or language models for textual descriptions [timemmd]. Through this approach, local context from multiple sources and domains can be seamlessly embedded in the framework to guide the revising process.
The hidden states are then passed through a traditional Transformer [vaswani2017attention] with a linear prediction head to generate the revised results $y_{local}$. By leveraging the attention mechanism, the local revision component explicitly captures the correlations between covariates and exogenous variables, thereby ensuring that the contextual information within the local window is fully utilized to correct forecasting failures and enhance prediction accuracy.
@@ 3.4 Global Revising
$\text{Index},w=$ $\text{TopK }\operatorname{Sim}(Enc(x),Enc(X)),$ (3)
$Y_{re}$ $=\{Y_{train,i}|\forall i\in\text{Index}\}.$
Here, $Enc(\cdot)$ is an encoding function that processes the time series, which is instantiated as an instance normalization operation [revin] in practice for its simplicity and effectiveness of mitigating the impact of non-stationarity that could lead to unstable similarity estimation. Moreover, powerful pre-trained forecasting models [timesfm, timer, moirai] can be optionally utilized for better representation projection. Additionally, $w$ represents the top-$K$ similarity scores estimated by the similarity operator $\operatorname{Sim}(\cdot,\cdot)$, which is instantiated using cosine similarity due to its computational efficiency.
@@ 3.5 Optimization Target
By combining the components described above, the PIR framework produces the final forecasting results through post-forecasting identification and revision using a residual approach [resnet]:
@@
Here, $\alpha$ and $\beta$ are learned weights corresponding to the local and global revision components, respectively, and $\sigma(\cdot)$ denotes the Sigmoid activation function. A linear transformation is employed to estimate $\alpha$, with its weight and bias initialized to vectors of ones and zeros. This design ensures ensuring a positive correlation that higher uncertainty estimates lead to larger values of $\alpha$, thus placing greater emphasis on the local revision. In contrast, $\beta$ is generated through a multi-layer perceptron (MLP), which takes both the estimated uncertainty and the retrieval similarities as input. This allows the model to dynamically adjust the influence of the global revision based on the confidence of the prediction and the relevance of the retrieved historical series. The overall optimization objective is to minimize the MSE between the revised forecasts and the ground truth, formulated as:
@@ 4.1 Experimental Setup
Datasets. For the long-term forecasting task, we conduct experiments on a widely recognized benchmark dataset that includes eight real-world datasets spanning diverse domains [wu2021autoformer, lstnet]. Additionally, we incorporate the PEMS dataset, which contains four subsets, for the short-term forecasting task [scinet]. The exogenous information used in these datasets are the available timestamps. We also conduct experiments on datasets with additional textual descriptions in the Appendix. Following standard experimental protocols, we split each dataset into training, validation, and testing sets in chronological order. The split ratios are set to 6:2:2 for the ETT dataset and 7:1:2 for the other datasets. Detailed information about the datasets is available in the Appendix.
Backbone Models. PIR is a model-agnostic plugin that can be seamlessly integrated with any time series forecasting model. To demonstrate the effectiveness of the framework, we select four mainstream forecasting models based on diverse architectures, encompassing both channel-dependent and channel-independent assumptions, as backbones. These models are evaluated under both long-term and short-term forecasting settings: PatchTST [patchtst], SparseTSF [sparsetsf], iTransformer [itransformer] and TimeMixer [timemixer]. We implement these models following their official code.